# Breaking the Dimensional Barrier: A Pontryagin-Guided Direct Policy Optimization for Continuous-Time Multi-Asset Portfolio

### v3

@@ 2.2 Optimality Conditions via Pontryagin’s Maximum Principle
We now apply Pontryagin’s Maximum Principle (PMP) to the continuous-time portfolio problem formulated in Section 2.1, where market parameters depend on the exogenous state process $\{\mathbf{Y}_{t}\}_{0\leq t\leq T}$. PMP provides necessary conditions for optimality through a system of forward-backward stochastic differential equations (FBSDEs), which couple the dynamics of the state variables with those of associated costate variables.
@@
The system state includes the investor’s wealth $X_{t}$ and the exogenous state variables $\mathbf{Y}_{t}\in\mathbb{R}^{k}$. The corresponding costate processes represent the sensitivity of the value function $V$ to changes in the state: $\lambda_{t}=V_{x}(t,X_{t},\mathbf{Y}_{t})$ is the scalar costate for wealth (marginal value of wealth), and $\bm{\eta}_{t}=V_{\mathbf{y}}(t,X_{t},\mathbf{Y}_{t})$ is the $k$-dimensional costate vector for the exogenous state variables (marginal value of state variables).
Under an optimal control policy $\bigl{(}\bm{\pi}_{t}^{*},\,C_{t}^{*}\bigr{)}$, the evolution of the system state, comprising the wealth process $X_{t}^{*}$ and the exogenous variables $\mathbf{Y}_{t}$, is described by the following forward stochastic differential equations:
$dX_{t}^{*}$ $=\Bigl{[}X_{t}^{*}\,\Bigl{(}r\bigl{(}t,\mathbf{Y}_{t}\bigr{)}+\bm{\pi}_{t}^{*\top}\,\bigl{(}\bm{\mu}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}-r(t,\mathbf{Y}_{t})\mathbf{1}\bigr{)}\Bigr{)}-C_{t}^{*}\Bigr{]}\,dt$
$\quad+X_{t}^{*}\,\bm{\pi}_{t}^{*\top}\,\bm{\sigma}(t,\mathbf{Y}_{t})\,d\mathbf{W}_{t}^{X},$ (4)
$d\mathbf{Y}_{t}$ $=\bm{\mu}_{Y}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\,dt+\bm{\sigma}_{Y}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\,d\mathbf{W}_{t}^{Y},$ (5)
@@
with initial conditions $X_{0}^{*}=x_{0}>x_{\min}$ and $\mathbf{Y}_{0}=\mathbf{y}_{0}\in\mathbb{R}^{k}$. Recall the definitions of the volatility matrix $\bm{\sigma}(t,\mathbf{Y}_{t})$ (for assets) and the covariance structure $\bm{\Omega}$ of the combined Brownian motion $\mathbf{W}_{t}=(\mathbf{W}^{X\top}_{t},\mathbf{W}^{Y\top}_{t})^{\top}$.
Complementing these forward dynamics, the optimal costate processes $(\lambda_{t}^{*},\bm{\eta}_{t}^{*})$ satisfy a system of backward stochastic differential equations (BSDEs). These equations are driven by derivatives of the Hamiltonian $\mathcal{H}$ associated with the control problem:
$-d\lambda_{t}^{*}$ $=\mathcal{H}_{x}\bigl{(}t,X_{t}^{*},\mathbf{Y}_{t},\bm{\pi}_{t}^{*},C_{t}^{*},\lambda_{t}^{*},\bm{\eta}_{t}^{*},\mathbf{Z}_{t}\bigr{)}\,dt-\mathbf{Z}_{t}^{\lambda X}\,d\mathbf{W}_{t}^{X}-\mathbf{Z}_{t}^{\lambda Y}\,d\mathbf{W}_{t}^{Y},$ (6)
$-d\bm{\eta}_{t}^{*}$ $=\mathcal{H}_{\mathbf{y}}\bigl{(}t,X_{t}^{*},\mathbf{Y}_{t},\bm{\pi}_{t}^{*},C_{t}^{*},\lambda_{t}^{*},\bm{\eta}_{t}^{*},\mathbf{Z}_{t}\bigr{)}\,dt-\mathbf{Z}_{t}^{\eta X}\,d\mathbf{W}_{t}^{X}-\mathbf{Z}_{t}^{\eta Y}\,d\mathbf{W}_{t}^{Y}.$ (7)
Here, $\mathcal{H}_{x}=\partial\mathcal{H}/\partial x$ and $\mathcal{H}_{\mathbf{y}}=\nabla_{\mathbf{y}}\mathcal{H}$ represent the partial derivatives of the Hamiltonian. The terms $\mathbf{Z}_{t}^{\lambda X}\in\mathbb{R}^{1\times n}$, $\mathbf{Z}_{t}^{\lambda Y}\in\mathbb{R}^{1\times k}$, $\mathbf{Z}_{t}^{\eta X}\in\mathbb{R}^{k\times n}$, and $\mathbf{Z}_{t}^{\eta Y}\in\mathbb{R}^{k\times k}$ are adapted processes arising from the martingale representation theorem applied to the costates, collectively denoted by $\mathbf{Z}_{t}$. The derivatives $\mathcal{H}_{x},\mathcal{H}_{\mathbf{y}}$ generally depend on the state, costate, controls, and the $\mathbf{Z}_{t}$ terms themselves, reflecting the system’s interconnectedness through the value function’s second-order derivatives. The terminal conditions are determined by the objective functional: $\lambda_{T}^{*}=\kappa\,e^{-\delta T}\,U^{\prime}\bigl{(}X_{T}^{*}\bigr{)}$ and $\bm{\eta}_{T}^{*}=\mathbf{0}$.
The application of PMP involves maximizing a Hamiltonian function. To properly account for the dependence of the wealth diffusion on the control $\bm{\pi}$, the Hamiltonian should incorporate the interaction between the costate dynamics (represented by $\mathbf{Z}$ terms) and the diffusion coefficients. Following formulations in stochastic control theory (e.g., Yong and Zhou, , 2012), the appropriate Hamiltonian $\mathcal{H}$ can be expressed as:
@@
$\mathcal{H}\!=\!e^{-\delta t}U(C)+\lambda\!\left[x\bigl{(}r+\pi^{\top}(\mu-r\mathbf{1})\bigr{)}-C\right]+\bm{\eta}^{\top}\mu_{Y}+\mathbf{Z}^{\lambda X}\!\bigl{(}x\sigma^{\top}\pi\bigr{)}+\bigl{(}\mathbf{Z}^{\eta X}\bigr{)}^{\!\top}\!\bigl{(}x\sigma^{\top}\pi\bigr{)}+\bigl{(}\mathbf{Z}^{\eta Y}\bigr{)}^{\!\top}\sigma_{Y}.$ (8)
For the $\bm{\pi}$-FOC, only the drift term $\lambda x\bm{\pi}^{\top}(\bm{\mu}(t,\mathbf{Y}_{t})-r(t,\mathbf{Y}_{t})\mathbf{1})$ and the diffusion-coupling term $\mathbf{Z}^{\lambda X}(x\,\bm{\sigma}(t,\mathbf{Y}_{t})^{\top}\bm{\pi})$ matter for the maximization step.
@@
To ensure the well-posedness of this FBSDE system and the applicability of PMP, additional smoothness conditions on the model coefficients are typically required.
@@
    $C^{1}$-regular coefficients w.r.t. $\mathbf{y}$. All coefficient functions $\bm{\mu}_{Y},\,\bm{\sigma}_{Y},\,\bm{\mu},\,\bm{\sigma}\,$ $r$ are continuously differentiable with respect to the state vector $\mathbf{y}$. Furthermore, their partial derivatives ($\partial_{\mathbf{y}}f$) are globally Lipschitz continuous in $\mathbf{y}$ and satisfy linear growth conditions. This ensures that the drift terms in the costate BSDEs, involving derivatives like $\mathcal{H}_{\mathbf{y}}$, satisfy the required regularity conditions for the existence and uniqueness of solutions (cf. Yong and Zhou, , 2012).
The core requirement of Pontryagin’s Maximum Principle is that the optimal controls $(\bm{\pi}_{t}^{*},C_{t}^{*})$ must maximize the Hamiltonian $\mathcal{H}$ (8) pointwise in time, almost surely, given the optimal state and costate processes $(X_{t}^{*},\mathbf{Y}_{t},\lambda_{t}^{*},\bm{\eta}_{t}^{*},\mathbf{Z}_{t}^{*})$. Assuming interior solutions allows us to derive the first-order conditions (FOCs) for optimality. Maximization with respect to the consumption rate $C_{t}$ yields the standard condition:
@@
This classic result links optimal consumption $C_{t}^{*}=(U^{\prime})^{-1}(e^{\delta t}\lambda_{t}^{*})$ inversely to the marginal value of wealth $\lambda_{t}^{*}$. Maximization with respect to the portfolio weights $\bm{\pi}_{t}$ requires differentiating $\mathcal{H}$ with respect to $\bm{\pi}$, including the term involving $\mathbf{Z}^{\lambda X}$. From (8), the relevant terms are $\lambda x\bm{\pi}^{\top}(\bm{\mu}-r\mathbf{1})$ and $\mathbf{Z}^{\lambda X}(x\,\bm{\sigma}^{\top}\bm{\pi})$. Setting the gradient $\nabla_{\bm{\pi}}\mathcal{H}=\mathbf{0}$ gives (assuming $X_{t}^{*}>0$):
$\lambda_{t}^{*}X_{t}^{*}(\bm{\mu}(t,\mathbf{Y}_{t})-r(t,\mathbf{Y}_{t})\mathbf{1})+X_{t}^{*}\bm{\sigma}(t,\mathbf{Y}_{t})(\mathbf{Z}_{t}^{\lambda X*})^{\top}=\mathbf{0}.$
Dividing by $X_{t}^{*}$, we obtain the FOC for the optimal portfolio:
@@
This first-order condition implicitly defines the optimal portfolio $\bm{\pi}_{t}^{*}$ through its dependence on the costate $\lambda_{t}^{*}$ and the martingale component $\mathbf{Z}_{t}^{\lambda X*}$.
To obtain an explicit expression for $\bm{\pi}_{t}^{*}$, we need to relate the martingale component $\mathbf{Z}^{\lambda X*}$ to the derivatives of the value function. Applying Ito’s lemma to $\lambda_{t}=V_{x}(t,X_{t},\mathbf{Y}_{t})$ and matching terms with the BSDE (6) establishes this relationship (the precise derivation often involves careful application of Ito’s formula for processes driven by correlated Brownian motions, see e.g., standard texts or appendices like Appendix A). This procedure yields an expression for $\mathbf{Z}^{\lambda X*}$ in terms of $V_{xx}$, $V_{x\mathbf{y}}$, the optimal controls $\bm{\pi}_{t}^{*}$, and the model parameters ($\bm{\sigma},\bm{\sigma}_{Y},\Psi,\rho$). Substituting this expression into the FOC (10) and solving the resulting algebraic equation for $\bm{\pi}_{t}^{*}$ leads to the well-known optimal portfolio formula:
$\bm{\pi}_{t}^{*}$ $=-\frac{V_{x}}{X_{t}V_{xx}}\,\Sigma^{-1}(t,\mathbf{Y}_{t})\,\bigl{(}\bm{\mu}(t,\mathbf{Y}_{t})-r(t,\mathbf{Y}_{t})\mathbf{1}\bigr{)}$ (11)
$\qquad\qquad-\frac{1}{X_{t}V_{xx}}\,\Sigma^{-1}(t,\mathbf{Y}_{t})\,\bigl{(}\bm{\sigma}(t,\mathbf{Y}_{t})\,\bm{\rho}\,\bm{\sigma}_{Y}(t,\mathbf{Y}_{t})\bigr{)}V_{x\mathbf{y}}.$
Here, $\Sigma(t,\mathbf{Y}_{t})=\bm{\sigma}(t,\mathbf{Y}_{t})\Psi\bm{\sigma}(t,\mathbf{Y}_{t})^{\top}$ represents the instantaneous covariance matrix of the risky asset returns. The value function derivatives $V_{x}=\lambda_{t}^{*}$, $V_{xx}=\partial_{x}\lambda_{t}^{*}$, and $V_{x\mathbf{y}}=\nabla_{\mathbf{y}}\lambda_{t}^{*}$ (a $k\times 1$ vector) are evaluated along the optimal path $(t,X_{t}^{*},\mathbf{Y}_{t})$. The formula clearly separates the portfolio into two components: the myopic demand, which optimizes the immediate risk-return trade-off as if the investment opportunity set were constant, and the intertemporal hedging demand, which adjusts the portfolio to mitigate risks arising from future stochastic changes in the state variables $\mathbf{Y}_{t}$. The hedging demand’s magnitude and direction depend on the correlation $\rho$ between asset returns and state variable innovations, and crucially on $V_{x\mathbf{y}}$, which measures how the marginal value of wealth changes with the state variables.
@@
For illustration, in the simplified case with one risky asset ($n=1$) and one state variable ($k=1$), the optimal allocation (11) becomes:
$\pi_{t}^{*}=-\frac{V_{x}}{X_{t}V_{xx}\sigma^{2}}\bigl{(}\mu(t,Y_{t})-r(t,Y_{t})\bigr{)}-\frac{1}{X_{t}V_{xx}\sigma^{2}}(\sigma\rho\sigma_{Y})V_{xy},$
clearly showing the separation between the myopic component, proportional to the market price of risk, and the hedging component, proportional to $V_{xy}$ and the correlation $\rho$.

### v4

@@ 2.2 Optimality Conditions via Pontryagin’s Maximum Principle
We now apply Pontryagin’s Maximum Principle (PMP) to the continuous-time portfolio choice problem formulated in Section 2.1, where market parameters depend on the exogenous state process $\{\mathbf{Y}_{t}\}_{0\leq t\leq T}$. PMP provides necessary conditions for optimality via a system of forward-backward stochastic differential equations (FBSDEs), which couple the dynamics of the state variables with those of the associated costate (or adjoint) processes.
@@
The system state consists of the investor’s wealth $X_{t}$ and the exogenous state variables $\mathbf{Y}_{t}\in\mathbb{R}^{k}$. The corresponding costate processes represent the marginal value of state variables with respect to the value function $V$: the scalar process $\lambda_{t}:=V_{x}(t,X_{t},\mathbf{Y}_{t})$ corresponds to the sensitivity to wealth, and the vector process $\bm{\eta}_{t}:=V_{\mathbf{y}}(t,X_{t},\mathbf{Y}_{t})\in\mathbb{R}^{k}$ captures sensitivity to the exogenous state.
Under an optimal control policy $(\bm{\pi}_{t}^{*},C_{t}^{*})$, the forward state dynamics are governed by:
$dX_{t}^{*}$ $=\left[X_{t}^{*}\,\bigl{(}r(t,\mathbf{Y}_{t})+\bm{\pi}_{t}^{*\top}\bigl{(}\bm{\mu}(t,\mathbf{Y}_{t})-r(t,\mathbf{Y}_{t})\mathbf{1}\bigr{)}\bigr{)}-C_{t}^{*}\right]dt+X_{t}^{*}\,\bm{\pi}_{t}^{*\top}\bm{\sigma}(t,\mathbf{Y}_{t})\,d\mathbf{W}_{t}^{X},$ (4)
$d\mathbf{Y}_{t}$ $=\bm{\mu}_{Y}(t,\mathbf{Y}_{t})\,dt+\bm{\sigma}_{Y}(t,\mathbf{Y}_{t})\,d\mathbf{W}_{t}^{Y},$ (5)
@@
with initial conditions $X_{0}^{*}=x_{0}>x_{\min}$ and $\mathbf{Y}_{0}=\mathbf{y}_{0}\in\mathbb{R}^{k}$. Recall that $\bm{\sigma}(t,\mathbf{Y}_{t})$ is the asset volatility matrix, and $\bm{\Omega}$ is the covariance matrix of the combined Brownian motion $\mathbf{W}_{t}=\bigl{(}\mathbf{W}^{X\top}_{t},\,\mathbf{W}^{Y\top}_{t}\bigr{)}^{\top}$.
To characterize optimal controls, PMP introduces a Hamiltonian function that incorporates both the instantaneous utility and the effects of the controls on the system’s drift and diffusion. In problems with control-dependent diffusion, the Hamiltonian must also account for the interaction with costate diffusion terms (i.e., the $\mathbf{Z}$-components of the adjoint BSDEs). Following standard stochastic control formulations (e.g., Yong and Zhou, , 2012), the Hamiltonian $\mathcal{H}$ is defined as:
@@
$\mathcal{H}=e^{-\delta t}\,U(C)+\lambda\left[x\bigl{(}r+\bm{\pi}^{\top}(\bm{\mu}-r\mathbf{1})\bigr{)}-C\right]+\bm{\eta}^{\top}\bm{\mu}_{Y}+\mathbf{Z}^{\lambda X}\cdot(x\,\bm{\sigma}^{\top}\bm{\pi})+(\mathbf{Z}^{\eta Y})^{\top}\bm{\sigma}_{Y},$ (6)
where $\lambda$ and $\bm{\eta}$ are the costate processes, $\mathbf{Z}^{\lambda X}$, $\mathbf{Z}^{\eta X}$, and $\mathbf{Z}^{\eta Y}$ represent the diffusion coefficients of the backward SDEs associated with $\lambda_{t}$ and $\bm{\eta}_{t}$, and all functions are evaluated at $(t,x,\mathbf{y},\bm{\pi},C)$, though we suppress this for clarity.
Complementing the forward dynamics, the optimal costate processes $(\lambda_{t}^{*},\bm{\eta}_{t}^{*})$ satisfy a system of backward stochastic differential equations (BSDEs), driven by the partial derivatives of the Hamiltonian $\mathcal{H}$ with respect to the state variables:
$-d\lambda_{t}^{*}$ $=\mathcal{H}_{x}\bigl{(}t,X_{t}^{*},\mathbf{Y}_{t},\bm{\pi}_{t}^{*},C_{t}^{*},\lambda_{t}^{*},\bm{\eta}_{t}^{*},\mathbf{Z}_{t}^{*}\bigr{)}\,dt-\mathbf{Z}_{t}^{\lambda X*}\,d\mathbf{W}_{t}^{X}-\mathbf{Z}_{t}^{\lambda Y*}\,d\mathbf{W}_{t}^{Y},$ (7)
$-d\bm{\eta}_{t}^{*}$ $=\mathcal{H}_{\mathbf{y}}\bigl{(}t,X_{t}^{*},\mathbf{Y}_{t},\bm{\pi}_{t}^{*},C_{t}^{*},\lambda_{t}^{*},\bm{\eta}_{t}^{*},\mathbf{Z}_{t}^{*}\bigr{)}\,dt-\mathbf{Z}_{t}^{\eta X*}\,d\mathbf{W}_{t}^{X}-\mathbf{Z}_{t}^{\eta Y*}\,d\mathbf{W}_{t}^{Y}.$ (8)
Here, $\mathcal{H}_{x}:=\partial\mathcal{H}/\partial x$ and $\mathcal{H}_{\mathbf{y}}:=\nabla_{\mathbf{y}}\mathcal{H}$ denote the partial derivatives of the Hamiltonian with respect to wealth and the state vector, respectively.
The terms $\mathbf{Z}_{t}^{\lambda X}\in\mathbb{R}^{1\times n}$, $\mathbf{Z}_{t}^{\lambda Y}\in\mathbb{R}^{1\times k}$, $\mathbf{Z}_{t}^{\eta X}\in\mathbb{R}^{k\times n}$, and $\mathbf{Z}_{t}^{\eta Y}\in\mathbb{R}^{k\times k}$ are adapted processes. [^7] These terms arise from the martingale representation theorem applied to the backward dynamics of the costate variables. See Karatzas and Shreve, (1991) for a detailed treatment of the martingale representation theorem. These processes are collectively denoted by $\mathbf{Z}_{t}$.
Note that $\mathcal{H}_{x}$ and $\mathcal{H}_{\mathbf{y}}$ generally depend not only on the state, costate, and control variables, but also on the $\mathbf{Z}_{t}$ terms themselves. This reflects the coupling between the forward and backward systems, which arises from the second-order sensitivity of the value function.
The terminal conditions for the BSDEs are determined by the structure of the objective functional:
$\lambda_{T}^{*}=\kappa\,e^{-\delta T}\,U^{\prime}\bigl{(}X_{T}^{*}\bigr{)},\qquad\bm{\eta}_{T}^{*}=\mathbf{0}.$
@@
To ensure the well-posedness of the forward–backward stochastic differential equation (FBSDE) system and the applicability of Pontryagin’s Maximum Principle (PMP), additional smoothness conditions on the model coefficients are typically required.
@@
    $C^{1}$-regularity with respect to $\mathbf{y}$. All coefficient functions—$\bm{\mu}_{Y},\,\bm{\sigma}_{Y},\,\bm{\mu},\,\bm{\sigma},\,r$—are continuously differentiable with respect to the state vector $\mathbf{y}$. Moreover, their partial derivatives $\partial_{\mathbf{y}}f$ are globally Lipschitz continuous in $\mathbf{y}$ and satisfy linear growth bounds. These conditions ensure that the drift terms in the costate BSDEs, including expressions such as $\mathcal{H}_{\mathbf{y}}$, satisfy the regularity requirements for the existence and uniqueness of solutions; see Yong and Zhou, (2012, Chapter 2 and 6).
The core requirement of Pontryagin’s Maximum Principle is that the optimal controls $(\bm{\pi}_{t}^{*},C_{t}^{*})$ maximize the Hamiltonian $\mathcal{H}$ (6) pointwise in time, almost surely, given the optimal state and costate processes $(X_{t}^{*},\mathbf{Y}_{t},\lambda_{t}^{*},\bm{\eta}_{t}^{*},\mathbf{Z}_{t}^{*})$. Assuming interior solutions allows us to derive the first-order conditions (FOCs) for optimality. Maximizing with respect to the consumption rate $C_{t}$ yields the standard FOC:
@@
This classical result links optimal consumption to the marginal value of wealth via the inverse of the marginal utility: $C_{t}^{*}=(U^{\prime})^{-1}(e^{\delta t}\lambda_{t}^{*})$.
For the portfolio weights $\bm{\pi}_{t}$, we differentiate $\mathcal{H}$ with respect to $\bm{\pi}$, accounting for both the drift and diffusion terms that depend on the control. Setting the gradient $\nabla_{\bm{\pi}}\mathcal{H}=\mathbf{0}$, we obtain:
@@
Dividing by $X_{t}^{*}$ assuming $X_{t}^{*}\neq 0$, the first-order condition for the optimal portfolio becomes:
[…]
@@
This condition implicitly defines the optimal portfolio $\bm{\pi}_{t}^{*}$ through its dependence on the costate $\lambda_{t}^{*}$ and the martingale component $\mathbf{Z}_{t}^{\lambda X*}$, which captures the diffusion sensitivity of the marginal value of wealth.
To obtain an explicit expression for $\bm{\pi}_{t}^{*}$, we must relate the coefficient $\mathbf{Z}^{\lambda X*}$ of the martingale component in the BSDE (7) to the derivatives of the value function. This relationship can be derived by applying Itô’s lemma to $\lambda_{t}=V_{x}(t,X_{t},\mathbf{Y}_{t})$ and matching terms with the dynamics of the costate process in (7). The full derivation is presented in Appendix A.
This procedure yields an expression for $\mathbf{Z}^{\lambda X*}$ in terms of the second-order derivatives $V_{xx}$, $V_{x\mathbf{y}}$, the optimal controls $\bm{\pi}_{t}^{*}$, and model parameters $(\bm{\sigma},\bm{\sigma}_{Y},\Psi,\rho)$. Substituting this into the first-order condition (10) and solving the resulting algebraic equation for $\bm{\pi}_{t}^{*}$ leads to the well-known optimal portfolio formula, originally derived by Merton, (1973):
[…]
$\quad-\frac{1}{X_{t}V_{xx}}\,\Sigma^{-1}(t,\mathbf{Y}_{t})\,\bm{\sigma}(t,\mathbf{Y}_{t})\,\bm{\rho}\,\bm{\sigma}_{Y}(t,\mathbf{Y}_{t})\,V_{x\mathbf{y}}.$
Here, $\Sigma(t,\mathbf{Y}_{t}):=\bm{\sigma}(t,\mathbf{Y}_{t})\,\Psi\,\bm{\sigma}(t,\mathbf{Y}_{t})^{\top}$ is the instantaneous covariance matrix of risky asset returns. The derivatives of the value function —$V_{x}=\lambda_{t}^{*}$, $V_{xx}=\partial_{x}\lambda_{t}^{*}$, and $V_{x\mathbf{y}}=\nabla_{\mathbf{y}}\lambda_{t}^{*}\in\mathbb{R}^{k}$—are evaluated along the optimal trajectory $(t,X_{t}^{*},\mathbf{Y}_{t})$.
This expression decomposes the optimal portfolio into two components: a myopic demand, which optimizes the static trade-off between expected return and risk under fixed investment opportunities, and an intertemporal hedging demand, which adjusts the portfolio to hedge against stochastic shifts in future investment opportunities. The hedging component is shaped by the correlation structure $\bm{\rho}$ between asset returns and state innovations, and by the cross-derivative $V_{x\mathbf{y}}$, which captures how the marginal value of wealth responds to changes in the economic state.
@@
For illustration, consider the simplified case with a single risky asset ($n=1$) and a single state variable ($k=1$). In this setting, the optimal portfolio allocation in (11) reduces to:
$\pi_{t}^{*}=-\frac{V_{x}}{X_{t}V_{xx}\sigma^{2}}\bigl{(}\mu(t,Y_{t})-r(t,Y_{t})\bigr{)}-\frac{1}{X_{t}V_{xx}\sigma^{2}}\,\sigma\rho\sigma_{Y}\,V_{xy},$
where $\sigma:=\sigma(t,Y_{t})$, $\sigma_{Y}:=\sigma_{Y}(t,Y_{t})$, and $V_{xy}:=\partial^{2}V/\partial x\partial y$.
This expression makes explicit the decomposition of the optimal policy into two components: a myopic demand, which is proportional to the market price of risk $(\mu-r)/\sigma$, and a hedging demand, which is proportional to the cross-partial derivative $V_{xy}$ and the correlation coefficient $\rho$ between shocks to the asset return and the state variable.